OpenStudy (anonymous):

I just wanted to know how to solve a system of equations when there is 3 equations. I don't want the answers I just want to know how to do it

OpenStudy (hba):

You will be suspended from school if you are trying to get the answers.

OpenStudy (hba):

@Deepcobra17 If you need help post your question one by one rather than writing connections algebra 2 students and stuff please.

OpenStudy (anonymous):

I just wanted to know how to solve a system of equations when there is 3 equations. I don't want the answers I just want to know how to do it

OpenStudy (hba):

@Deepcobra17 You would have simply posted that rather than saying [CONNECTIONS ACADEMY AND STUFF] Thanks a lot for understanding :)

OpenStudy (amistre64):

the same methods used to solve 2 equations are used to solve 3

OpenStudy (anonymous):

I asked my math teacher and he doesn't want to help so I figured other people from that school could explain it better then he could. But it's okay, I'll go watch a YouTube explaining video or something

OpenStudy (anonymous):

So I can use elimination?

OpenStudy (amistre64):

you can; use one equation to eliminate a variable in the other 2; then you can eliminate again using the modified equations

OpenStudy (amistre64):

a specific problem to work thru would help alot :)

OpenStudy (anonymous):

-3x - 4y -3z = -7 2x -6y + 2z = 3 5x - 2y + 5z = 9

OpenStudy (amistre64):

since the y parts all have a common factor; those look to be the simplest to eliminate to start with to me; what is your idea?

OpenStudy (amistre64):

-3x - 4y -3z = -7 2x -6y + 2z = 3 5x - 2y + 5z = 9 <-- use this to elimiate the ys with -15x + 6y -15z = -27 2x -6y + 2z = 3 ---------------------- -13x -13z = -24 -10x +4y - 10z = -18 - 3x - 4y - 3z = - 7 -------------------- -13x -13z = -25

OpenStudy (amistre64):

the modified results look to me like they have no common solution so the overall system is inconclusive

OpenStudy (anonymous):

Hold on, I'm still stuck on the first step

OpenStudy (amistre64):

tell me where you are stuck at

OpenStudy (anonymous):

How you got -15 + 6y

OpenStudy (amistre64):

notice the 3rd equation has a -2y in it; and the others have a 6 and a 4 ... i used multiples of the 3rd equation to eliminate the y parts in the other 2

OpenStudy (amistre64):

3rd eq: 5x - 2y + 5z = 9 times -3 to get -15x +6y -15z = -27 i multiplied it by -2 for the other one

OpenStudy (anonymous):

Oh okay, I think I understand now

OpenStudy (amistre64):

the modified results, if i did the mathing right :) appears to be parallel planes (or lines or whatever you want to call them) that never meet

OpenStudy (anonymous):

Ok, so would I then Plug in the values to solve for x, y, and z?

OpenStudy (anonymous):

Never mind I just worked it out on paper, thank you so much!

OpenStudy (amistre64):

youre welcome, and good luck. as i see it, there is no solution to the system